I need help with the rate constants for trials 1 through 4. I put as much info as I thought would be necessary. The last question I posted I was told I didn't have enough info, so I hope this is enough.

This is all the info I was given to solve the questions, so this has to be enough to solve them. Rate Constant, k, for the reaction occurring in Solution 1 (M1.51) NG 0.5 Info Rate Constant, k, for the reaction occurringTrial H2O2 (mL) KI (mL) H20 (ml) [H202] After Mixing (M) [Ki] After Mixing (M) Try 1: 70.700 Try 1: 70.101 4.1 Edit 1.05 EditThe test tubes used have a total volume of 22 ml. The gas will occupy 17 mL of the total volume. Trial Temperature (°C) It isTrial Temperature (°C) It is OKAY if the data looks out of line. Rate formation 02 (torr/s) Rate formation O2 (mol/Los-1) Ratne (5) Temperature Trial it is OKAY if the data looks out of line. Rate formation O2 (torr/s) Rate formation O2 (mol/Los-1) RTrial Temperature (°C) It is OKAY if the data looks out of line. Rate formation O2 (torr/s) Rate formation O2 (mol/Les 1) Rat
Answer #1

| H₂O KI after I KI Trial R02 mi H₂O2 after mixing ml ml CM7 1:05 0.700 0.101 3.99 0.5 0.5 0.703 0.050 |-00 2:0 0.362 0.098 2

Date Rate of Reaction = -1 d[ Hz03] British Rale el Reaction I Trial 1 R = 9.9E-005 - 000099 22 - 6-9503 Trial2 R2 = 0000 256Date: Now Calculation for Rate constant Initial Rate = K (H₂O₂] [I] Initial Rate 4.95e-5 [H027 - I - 0.700 0.101 Trial 2 K= I